OpenStudy (anonymous):

square root of t-8 = 2t+1

OpenStudy (anonymous):

t=-9

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you are going to get complex solutions

OpenStudy (anonymous):

using quadratic formula on 4t^2+3t+9

OpenStudy (anonymous):

t-8=(2t+1)^2 t-8=4t^2+4t+1 4t^2+3t+9=0 no real solution

OpenStudy (anonymous):

there is but they are complex imaginary

OpenStudy (anonymous):

can be that how do i get an answer

OpenStudy (anonymous):

using quadratic, do you remember that?

OpenStudy (anonymous):

yes but getting it wrong

OpenStudy (anonymous):

what are you getting

OpenStudy (anonymous):

weird answers

OpenStudy (anonymous):

3x3